OpenStudy (anonymous):

give the exact and approximate solutions to three decimal places (x-3)^2 =16

OpenStudy (anonymous):

\[x-3=\sqrt{16}\]x=4+3 x=7

OpenStudy (anonymous):

it says 7 is one answer its asking for two im confused

OpenStudy (anonymous):

they always ask for two..

OpenStudy (anonymous):

I for got \[x-3=\pm \sqrt{16}\]\[x=3\pm4\]

OpenStudy (anonymous):

they do?

OpenStudy (anonymous):

x=7 or x= -1

OpenStudy (anonymous):

ok i think

OpenStudy (anonymous):

ill post another one can you shw me step by step so i can try